OpenStudy (bigmoee19):

The vertex of this parabola is at (-4,-1). When the y-value is 0, the x-value is 2. What is the coefficient of the squares term in the parabola equation A:3 B:-6 C:-3 D:6

OpenStudy (bigmoee19):

Help please👍

OpenStudy (scarlettfarra2000):

Remember that the vertex form is: y = a(x - h)² + k

OpenStudy (zzr0ck3r):

\(y=a(x-h)^2+k\implies y=a(x-(-4))^2-1\\ \implies y=a(x+4)^2-1\) We use the given point \((0,2)\) to solve for \(a\). So let \(y=2\) and \(x=0\) in the above equation \(2=a(0+4)^2-1\) Can you solve for \(a\)?

OpenStudy (bigmoee19):

How will i plug them in?

OpenStudy (zzr0ck3r):

I did that part. We have the equation \(y=a(x+4)^2-1\) and we were given the point \((0,2)\), which means the equation is true when \(x=0\) and \(y=2\) So i plug in \(x=0\) and \(y=2\) and I get \(2=a(0+4)^2-1\) Notice the difference in \(y=(x+4)^2-1\) and \(2=a(0+4)^2-1\).

OpenStudy (bigmoee19):

Yes cause the equation for going left or right is with X= NOT Y=

OpenStudy (bigmoee19):

But i will end up getting 2=a (4)^2-1

OpenStudy (zzr0ck3r):

solve for \(a\)

OpenStudy (zzr0ck3r):

add \(1\) to both sides, and then divide both sides by \(16\) and you are done.

OpenStudy (bigmoee19):

But it wont get the anwser from above i get 5.33

OpenStudy (zzr0ck3r):

oh sorry, I thought it said \(x=0\) and \(y=2\) but it is the opposite. So solve this \(0=a(2+4)^2-1\) But I still don't see the right answer in the list :(